OpenStudy (anonymous):

limt xto0 sqrt1-x -sqrt1-x/x

OpenStudy (anonymous):

can you try entering this with the use of \(\LaTeX\) ? It is really hard to understand what this question asks for \[\Large \lim_{x \to 0} \sqrt{1-x}- \frac{\sqrt{1-x}}{x}? \]

OpenStudy (anonymous):

the x is for total not for sqrt 1-x

OpenStudy (anonymous):

please tell me the answer

OpenStudy (anonymous):

so it is \[\Large \lim_{x \to 0 } \frac{\sqrt{1-x}-\sqrt{1-x}}{x}? \]

OpenStudy (anonymous):

we need to agree on that first :-) \[\Large \lim_{x \to 0} \sqrt{1-x} - \sqrt{\frac{1-x}{x}} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the last one I posted?

OpenStudy (anonymous):

@venkatesh87 , if it's the last one I posted try to expand it with another root so you get the binomial form \((a+b)(a-b)=a^2-b^2\) if you need assistance with that, you can tell me. However I would first need clarification if we're talking about the same problem.